OpenStudy (anonymous):

What is the value of the fourth term in a geometric sequence for which a1=10 and r=0.5?

OpenStudy (jdoe0001):

\(\large { a_1\cdot r^{{\color{brown}{ n}}-1}\qquad \textit{so the 4th term is} \begin{array}{ccllll} term&value \\\hline\\ {\color{brown}{ 1}}&a_1\cdot r^{{\color{brown}{ 1}}-1}\\ {\color{brown}{ 2}}&a_1\cdot r^{{\color{brown}{ 2}}-1}\\ {\color{brown}{ 3}}&a_1\cdot r^{{\color{brown}{ 3}}-1}\\ ...&... \end{array} }\)